the water flows in the channel 844 mm x 586 mm at the rate of 8 m/sec. What is the volume flow in ft^3/sec?


1
Expert's answer
2021-12-06T09:45:00-0500

Given:

"A=\rm 844\: mm\times 586\: mm=0.495\: m^2"

"v=\rm 8 \: m/s"


The volume flow

"\frac{dV}{dt}=Av\\
=\rm 0.495\: m^2*8\: m/s=3.96\: m^3/s=140\: ft^3/s"